Circular No. 7568

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001M IN NGC 3240 A. B. Aazami and W. D. Li, University of California at Berkeley, on

behalf of LOTOSS (cf. IAUC 7514), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Jan. 21.4 (mag ∼ 17.5) and 22.4 UT (mag ∼ 17.7). SN 2001M is located at α = 10^{h}24^{m}31^{s}.48, δ = −21^{o}47′40′′.6 (equinox 2000.0), which is 12′′.3 east and 12′′.7 south of the nucleus of NGC 3240. A KAIT image of the same field taken on Jan. 3.4 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001N IN NGC 3327 Li also reports the discovery by R. Chornock of an apparent supernova

on unfiltered KAIT images taken on Jan. 21.4 (mag ∼ 16.3) and 22.4 UT (mag ∼ 16.2). SN 2001N is located at α = 10^{h}39^{m}58^{s}.08, δ = +24^{o}05′26′′.0 (equinox 2000.0), which is 1′′.8 east and 2′′.7 south of the nucleus of NGC 3327. The photometry and astrometry of the new object were measured after subtracting a KAIT image of the same field taken on Jan. 1.4, which showed nothing at the position of the new object (limiting mag ∼ 19.0).

MARKARIAN 421 H. G. Boerst, University of Kiel; N. Goetting, University of Hamburg;

and R. Remillard, Massachusetts Institute of Technology, report on behalf of the High-Energy Gamma-Ray Astronomy (HEGRA) collaboration and the ASM/RXTE team: “Following the detection of enhanced x-ray emission from the BL Lac object Markarian 421 with the All-Sky Monitor onboard the RXTE satellite, observations with the stereoscopic Cherenkov telescope system and the Cherenkov telescope CT1 of HEGRA at La Palma performed on Jan. 16–22 showed very strong TeV activity from the source. Above a threshold energy of 500 GeV, diurnal integral fluxes (20-min averages) of up to 1, 4, 2, 1.5, 2.5, and 3 times that of the Crab nebula have been measured during these consecutive nights (source at altitudes > 45^{o}). Observations at other wavelengths are strongly encouraged. For updates on the activity of Mkn 421 and the HEGRA observation schedule,

see http://www-hegra.desy.de/mrk-421/.”

2001 January 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green